whats is the mass in grams of 5.27x10^22 atoms of tin
1
Expert's answer
2016-01-29T09:40:29-0500
N(Sn)=5.27x10^22/6.02x10^23= 0.0875 mol
m(Sn)= 0.0875*118.69= 10.39 g